# Visual Enhanced Depth Scaling for Multimodal Latent Reasoning

### 16 Apr 2026 v2

@@ 4.3 Ablation Analysis
Table 5: Ablation Experiments. We provide ablation analysis of key parameters and experimental settings on three benchmarks: M3CoT, ScienceQA, and GQA. All the variants adopt Qwen2.5-VL-3B as the base model. Base refers the model that adopts RDS module and SCF-VR module.
@@
Visual Latents Formation.
@@ 4.4 In-depth Analysis
Impact of $\lambda$.
As shown in Figure 4, model performance exhibits high sensitivity to variations in the hyperparameter $\lambda$. Notably, larger models (e.g., 7B) achieve peak performance at $\lambda=1.0$, whereas the 2B model attains optimal results at a substantially lower value of $\lambda=0.2$. We attribute this divergence to distinct capacity across model scales: smaller models are primarily bottlenecked by visual perception capabilities, thus requiring larger degree of visual play to establish effective visual grounding for downstream reasoning. In contrast, larger models possess sufficient grounding capability, and they benefit from a more balanced allocation between visual and linguistic signals, avoiding over-reliance on visual features that could otherwise suppress the development of complex reasoning chains.
Figure 4: Sensitivity analysis of $\lambda$.
Figure 5: (a) and (b) illustrate the performance with different number of event prototypes and different ratio of filtered event prototypes, respectively.
Impact of different window size $W$.
Figure 6: Visualization of cropped region in each latent reasoning step. Dotted line denotes the correspondence between reasoning rationales and cropped region.
@@
As depicted in Figure 5, we visualize the embedding space distribution of multimodal features for both the baseline and our method. The latent tokens learned by our approach form several distinct clusters, situated centrally within the text embedding manifold. Notably, compared to the baseline, our latents exhibit closer proximity to visual embeddings. This observation suggests that our latent tokens encapsulate richer reasoning semantics while facilitating deeper integration of visual information.
Table 6: Performance comparison of different sub-grid window sizes on the M3CoT benchmark. We set $\lambda=1.0$ for the 2B/3B models and $\lambda=0.2$ for the 7B model. Here, $W$ denotes the sub-grid window size, with $W=10$ is the window size of extracted visual feature.
M3CoT (Acc.%)
Method Grid Size $W$ = 2 $W$ = 3 $W$ = 5
Qwen2-VL-2B 10 $\times$ 10 50.8 51.2 50.3
Qwen2-VL-7B 65.7 66.4 65.2
Qwen2.5-VL-3B 65.0 65.3 64.5
Qwen2.5-VL-7B 73.4 73.9 72.1
Visualization of Crop Region.
As illustrated in Figure 6, we visualize the attention-guided crops across successive reasoning steps. Taking the third case as a representative example, which queries the purpose of the train cart, the model initially localizes the region corresponding to the description “shape like a house” during the first reasoning step. In subsequent steps, the attention focus progressively shifts toward the “entrance door”. Remarkably, without any explicit fine-grained supervision, our curriculum training paradigm enables latent tokens to progressively capture logical dependencies and establish robust visual-semantic alignments. This empirically demonstrates the model’s capacity to jointly perform multi-step visual grounding and generate coherent reasoning chains.
Impact of Position Encoding.
Regarding positional encoding during depth scaling, we evaluate two strategies. The Keep Old Position variant preserves the original relative positional embeddings of the selected tokens. In contrast, the Rearrange Position variant sequentially re-indexes the selected tokens from $1$ to $K$ according to the reduced sequence length. As demonstrated in Table 3, preserving the original positions yields superior performance. Although re-indexing may enhance local contextual modeling among salient tokens, it risks introducing positional inconsistencies with embeddings from preceding reasoning steps and discards inherent structural priors, ultimately leading to suboptimal results.

### 11 May 2026 v4

@@ 3.4 Training Procedure
4 Experiments
4.1 Experimental Setup
Training Dataset and Evaluation Benchmark.
To facilitate reproducibility, we first detail the training data configuration. During the supervised fine-tuning (SFT) stage, we curate a subset of approximately 30K samples from OneThinker [10], selected for its diverse distribution of reasoning chain lengths. The statistical distributions of CoT length, reasoning steps, and topic categories within this subset are illustrated in Fig. 3. We evaluate our method on three tasks across twelve benchmarks: (1) Mathematics Reasoning (MathVista [36], MathVision [56], MM-Math [50]); (2) Vision-centric Reasoning (Hallusion-Bench [14], MMVP [53], Seed-Bench-2-Plus [30], HR-Bench [61], GQA [1]); (3) Multimodal Composition Reasoning (MMStar [4], BLINK [11], ScienceQA [37], M3CoT [5]). Additional descriptive details can be found in the Appendix.
Implementation Details.
To comprehensively evaluate the effectiveness and scalability of our approach, we instantiate our method across a diverse set of backbone architectures, including Qwen2-VL-2B/7B [57], Qwen2.5-VL-3B/7B [2], and Chameleon-7B [52].
We employ DeepSpeed ZeRO-2 optimization without CPU offloading, with a per-GPU batch size of 8. Training utilizes the Adam optimizer with a learning rate of $4\times 10^{-5}$ and $\beta_{1}=0.9$. For the proposed self-distillation loss, we set the weighting coefficient to $\lambda=1.0$ for 2B/3B-scale models and $\lambda=0.2$ for 7B-scale models, respectively. Input images are resized to 512 $\times$ 512 resolution when training on OneThinker subset and GQA, and resized to 280 $\times$ 280 on M3CoT and ScienceQA. All experiments are conducted on 16 NVIDIA H20 GPUs (96GB VRAM each). A single training run takes approximately 20 hours.
(a) CoT Length Distribution
(b) Thinking Step Distribution
(c) Topic Distribution
[…]
Baselines.
We evaluate our method against a comprehensive set of state-of-the-art baselines, which can be categorized into four paradigms: (1) Zero-shot VLMs (GPT-4o [41], LLaVA-OneVision [29], InternVL3.5-8B [59], Qwen2.5-VL-7B [2]), (2) Explicit CoT-based methods (SCAFFOLD [26], ICoT [13], Multimodal-CoT [80], CCoT [39], Chain-of-Focus [79]), (3) Visual Enhanced methods, including tool-augmented reasoning (DeepEyes [83]) and RL-enhanced VLM reasoning (Vision-R1 [24], PAPO [63], VL-Rethinker [55]), and (4) Multimodal Latent Reasoning approaches (Laser [62], LVR [27], Monet [58], DMRL [33]).
4.2 Overall Quantitative Results
Reasoning Accuracy.
As summarized in Table 1, trained with identical protocols, our method achieves the best results on three fine-grained CoT datasets across all backbones. Compared with explicit CoT baselines, it brings an average accuracy gain of nearly 30% on Qwen-2-VL. It also improves over the strongest latent baseline (IVT-LR) by an average of +2.63% across all benchmarks. As shown in Table 2, we further evaluate on mainstream diagnostic benchmarks. We observe consistent improvements on most tasks, with the largest gain on MMVP (+4.67%), which requires fine-grained visual discrimination. These improvements are attributed to our RCF-VR, which enforces spatially coherent visual latents to reduce hallucination while preserving visual details. In contrast, gains are more modest on text-intensive benchmarks (e.g., SeedBench-2-Plus), likely due to limited domain-specific data. Notably, despite operating purely in the latent space without external retrieval or RL-based optimization, our approach achieves competitive performance against explicit CoT and tool-augmented methods through token-wise depth scaling.
Reasoning Efficiency.
[…]
4.3 Ablation Analysis
Effect of Designed SCF-VR and RDS.
@@ 4.3 Ablation Analysis
4.4 Parametric Investigation
Table 5: Performance comparison of different sub-grid window sizes on the M3CoT benchmark. Here, $W$ denotes the sub-grid window size; $W=10$ is the window size of extracted visual features. We set $\lambda=1.0$ for the 2B/3B models and $\lambda=0.2$ for the 7B model.
M3CoT (Acc.%)
Method Grid Size $W=2$ $W=3$ $W=5$
Qwen2-VL-2B $\times10$ 50.8 51.2 50.3
Qwen2-VL-7B 65.7 66.4 65.2
Qwen2.5-VL-3B 65.0 65.3 64.5
Qwen2.5-VL-7B 73.4 73.9 72.1
Visual Latents Formation.
Impact of $\lambda$.
As shown in Fig. 4, model performance exhibits high sensitivity to variations in the hyperparameter $\lambda$. Notably, larger models (e.g., 7B) achieve peak performance at $\lambda=1.0$, whereas the 2B model attains optimal results at a substantially lower value of $\lambda=0.2$. We attribute this divergence to distinct capacity across model scales: smaller models are primarily bottlenecked by visual perception capabilities, thus requiring larger degree of visual play to establish effective visual grounding for downstream reasoning. In contrast, larger models possess sufficient grounding capability, and they benefit from a more balanced allocation between visual and linguistic signals, avoiding over-reliance on visual features that could otherwise suppress the development of complex reasoning chains.
Figure 4: Sensitivity analysis of $\lambda$.
Impact of different window size $W$.
@@ 4.4 Parametric Investigation
Figure 5: Visualization of learned router scores and selected top-$K$ tokens for depth scaling, revealing which textual tokens, CoT latent tokens, and visual regions exhibit greater optimization difficulty.
@@ 4.5 In-depth Analysis
Routing Behavior.
As illustrated in Fig. 5, the left panel highlights tokens selected by the learnable router (marked in red color). In our latent space reasoning framework, each reasoning step is compressed into a single latent token. The right panel displays the router scores for these latent tokens (circled in red), where the first and third latent token exhibit significantly higher scores, identifying them as difficult tokens requiring additional depth scaling. Specifically, the first reasoning step analyzes critical visual details essential for the answer, such as the skier’s arms touching the ground for balance, which indicates the difficulty level of the ski slope. Furthermore, the visualized attention scores in image reveal that the router prioritizes regions corresponding to snow, body limbs, and trees. This demonstrates that the model achieves deeper compositional reasoning by iteratively refining these semantic hubs.
Visualization of Crop Region.
As illustrated in Fig. 6, we visualize the attention-guided crops across successive reasoning steps. Taking the right-upper case as a representative example, which queries the purpose of the train cart, the model initially localizes the region corresponding to the description “shape like a house” during the first reasoning step. In subsequent steps, the attention focus progressively shifts toward the “entrance door”. Remarkably, without any explicit fine-grained supervision, our curriculum training paradigm enables latent tokens to progressively capture logical dependencies and establish robust visual-semantic alignments. This empirically demonstrates the model’s capacity to jointly perform multi-step visual grounding and generate coherent reasoning chains.
Figure 6: Visualization of cropped region in each latent reasoning step.
(a) Without RDS
(b) With RDS
Figure 7: (a) and (b) illustrate the Nuclear norm variance without RDS and with RDS.
(c) Vanilla
(d) Ours
Figure 8: (a) and (b) illustrate the t-SNE distribution of the learned latents, visual tokens, and text tokens.
Nuclear Norm Behavior with RDS.
As illustrated in Fig.87(a) and Fig.87(b), we conducted training experiments on a subset of hard samples. The results demonstrate that after incorporating the RDS module, the cumulative nuclear norm of the QKVO projection weights achieves stable convergence by the 3rd epoch. This observation suggests that the depth scaling strategy effectively mitigates gradient volatility and uncertainty induced by hard samples, thereby substantially enhancing the model’s capacity to learn and fit challenging instances and critical difficult tokens.
@@
As depicted in Fig. 87(c) and Fig. 87(d), we visualize the embedding space distribution of multimodal features for both the baseline and our method. The latent tokens learned by our approach form several distinct clusters, situated centrally within the text embedding manifold. Notably, compared to the baseline, our latents exhibit closer proximity to visual embeddings. This observation suggests that our latent tokens encapsulate richer reasoning semantics while facilitating deeper integration of visual information.